What is the answer to: n2 (n squared) - (minus) n=182 And how do i do it please?
Answer by scott8148(6628)

subtracting 182 ___ n^2 - n - 182 = 0

factoring ___ two factors of -182 that can "combine" to give -1 , like 13 and -14 ___ (n + 13)(n - 14) = 0

n + 13 = 0 ___ n = -13

n - 14 = 0 ___ n = 14